Shift summary for owl 8/1/99

Jason Hamm (jhamm@physics.arizona.edu)
Sun, 1 Aug 1999 05:05:16 -0700 (MST)


Shift Summary for owl shift, 08/1/99

On Shift: H. Nguyen, J. Hamm

Hours of beam available at all intensities: 5.7
Hours of beam available at nominal intensity: ~5.7
Hours of E832 data (special runs not included): 5.7
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: TeV problems
========================================================================
Runs ended this shift: none; 14077 (beam832) still running at end

Tapes brought to FCC this shift: QK*198-200, QKB068
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: 832 data taking at ~3.7 E12, with some interruptions
from TeV failures
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14077 0:00 Beam returns right as we take over.

14077 0:10 Beam gone: TeV lead failure, est. 10 minutes

14077 416 0:50 Beam returns, although spill structure is
initially poor. Structure and SEM return to
nominal after a few spills.

14077 1:25 Beam gone: Lost meson primary, no est.

14077 1:40 Beam returns

14077 522 3:15 Spill structure is awful. Following spills abort
about 1 second into T5. Problem due to bad TeV
corrector power supply, est. 30 min to replace.

14077 542 3:42 Beam returns

14077 4:30 Onsplit requests QKA005 to be mounted on ktev3,
we do so.

14077 590 4:50 Missing FFFF word in stream 5; kfrend fixes.

========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

CSI 1497 has ped rms of 3.25 over most of night. It has been put into
the monitoring namelist for the next run.